OpenStudy (anonymous):

Express the following functions as a positive angle: Cot (-99) Cos (-140)

OpenStudy (anonymous):

\[\cot (-99 {}^{\circ})=\tan (9 {}^{\circ}) \]\[\cos (-140 {}^{\circ})=-\cos (40 {}^{\circ}) \]

OpenStudy (anonymous):

\[\cot(-99)=\cos(-99)/\sin(-99)=\cos(99)/-\sin(99) = -\cot(99) \]